this question on my homework i dont get:
the sum of one number and thrice another is -15.the difference of the fisrt number and the second number is 10. find the numbers

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

the sum of one number and thrice another is -15.the difference of the fisrt number and the second number is 10. find the numbers
---------
Equation:
x + 3y = -15
x - y = 10
------------
Subtract and solve for "y":
4y = -25
y = 6 1/4
----------------
Solve for "x":
x - 6 1/4 = 10
-----
x = 3 3/4
================
Cheers,
Stan H.

Answer by rfer(16322)

x-y=10 x=10+y
--------------
x+3y=-15
10+y+3y=-15
4y=-25
y=-25/4
y=-6.25
-------
x+3*-6.25=-15
x-18.75=-15
x=-15+18.75
x=3.75
--------------
3.75+3*-6.25=-15
3.75-18.75=-15
-15=-15